OpenStudy (anonymous):

The graph below represents which system of inequalities? y > 2x - 3 y > -x - 3 y < 2x - 2 y < -x + 3 y 2x - 2 y > -x + 3 None of the above

OpenStudy (anonymous):

attach the graph

OpenStudy (rosedewittbukater):

You need to post the graph for us to help.

OpenStudy (anonymous):

^

OpenStudy (anonymous):

sorry, i was going to but got distracted

OpenStudy (rosedewittbukater):

Well the y intercept (where the line crosses the vertical axis) of the dash line is 3. A solid line means the sign is \[\le or \ge \] For the first line, it has to be y > mx + 3 or y < mx + 3. To find the slope m, find 2 points of the line. We can use (0, 3) and (1, 2). Subtract the first y from the second y and divide that by the first x subtracted from the second x. So 2-3/1-0. That makes -1. So the slope is -1. To find what the sign will be, replace all of the variables with 0. We'll use the = sign for now. y = -1x + 3 0 = -1(0) + 3 0 = 3 Since 0 is less than three, the sign is < The final equation for the dashed line is y < -x + 3 Can you find the equation for the solid line?

OpenStudy (anonymous):

well i think its C

OpenStudy (rosedewittbukater):

C has y > -x + 3. If you change the x and y to zero, you get 0 > 3. Is zero greater than 3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

....

OpenStudy (rosedewittbukater):

So it should be B

OpenStudy (anonymous):

no its not B

OpenStudy (anonymous):

@rosedewittbukater

OpenStudy (rosedewittbukater):

Why do you think that?

OpenStudy (rosedewittbukater):

The y intercept for the solid line is -2 and the slope is 2, meaning each point is 2 units up and 1 unit to the right of the point before it. y = 2x - 2 0 = -2 Zero is greater than negative 2, so the equation is y > 2x - 2 However, this is when the shading comes into play, which I forgot to mention before. We assume that the yellow shading is for the solid line only, and the green shading is where the shaded area of both of the lines overlap. Since the shading for the solid line is on the side that does not include the 0, the equation must be changed to y < 2x -2. As for the dashed line y < -x + 3 0 = 3 Since the shaded part is on the side of the line not including the 0, we switch the sign. The equation becomes y > -x +3. Sorry for my mistake earlier, I forgot that the picture included shading. So the answer is C.

OpenStudy (rosedewittbukater):

By the way, this is the english section for questions. This should be in math.